OpenStudy (anonymous):

Can someone please help me simplify this expression and put it in scientific notation? please! (4.5*10^-3)*512/(2.4*10^7)*(1.6*10^4)

OpenStudy (anonymous):

\[\frac{4.5*10^{-3}*512}{(2.4*10^7)*(1.6*10^4)}\]?

OpenStudy (anonymous):

Ya with the (4.5*10^-3) in ()

OpenStudy (anonymous):

ok, we dont need those (0 because multiplication is associative. but \[\frac{4.5*10^{-3}*512}{(2.4*10^7)*(1.6*10^4)}=\frac{4.5*512}{2.4*1.6}\frac{10^{-3}}{10^7*10^4}\\=\frac{4.5*512}{2.4*1.6}\frac{10^{-3}}{10^{7+4}}=\frac{4.5*512}{2.4*1.6} 10^{-3-11}=\frac{4.5*512}{2.4*1.6} 10^{-14}\]

OpenStudy (anonymous):

\[=600*10^{-14}=6*10^2*10^{-14} = 6*10^{2+(-14)}=6*10^{-12}\]

OpenStudy (anonymous):

if you think that 10^-14 is the answer, then you dont understand something. what is that something?

OpenStudy (anonymous):

even still I wrote = 4.58512 / 2.4*1.6 * 10^(-14) and you asked if it was 10^(-14) which makes it seem like you dont understand the steps...